OpenStudy (anonymous):

Find the surface area of a sphere with a diameter of 76 cm.

Parth (parthkohli):

Surface area of a sphere is: \(\Large \color{Black}{\Rightarrow ({d \over 2})^2 \times 4 \times \pi }\)

OpenStudy (anonymous):

thank ya !